# Isomorphism classes of generating sets

### v1

@@ 4. At a measurable cardinal
Let $a=(\alpha_{0},\ldots\alpha_{n-1},\beta)$ and and $a^{\prime}=(\alpha^{\prime}_{0},\ldots\alpha^{\prime}_{n-1},\beta^{\prime})$. If $\beta=\beta^{\prime}$, then $B_{a}\in\mathcal{X}_{a^{\prime}}$ and we have already seen that $\mathcal{X}_{a^{\prime}}\subseteq F_{a^{\prime}}$. Assume that $\beta^{\prime}<\beta$, then clearly $B_{a}\in V[G_{(\alpha_{1}+1)\times...\times(\alpha_{n-1}+1)\times\beta^{\prime}}]$. Let $\dot{B}_{a}$ be a canonical ${\mathbb{P}}_{a^{\prime}}$-name for $B_{a}$. We need to show that $\Vdash^{V}_{{\mathbb{P}}_{\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta^{\prime}}}\dot{B}_{a}\in A_{\beta^{\prime}}$. Let $G_{\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta^{\prime}}$ be an arbitrary generic for ${\mathbb{P}}_{\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta^{\prime}}$, and force to prolong it to a ${\mathbb{P}}$-generic $G$. Then by Claim 4.3
$B_{a}\in U\cap V[G_{\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta^{\prime}}]=U_{\beta^{\prime}}[G_{\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta^{\prime}}]=A_{\beta^{\prime}}[G_{\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta^{\prime}}]$
Hence, by definition, $\dot{B}_{a}\in F^{0}_{a^{\prime}}$ and $B_{a}\in F^{1}_{a^{\prime}}\subseteq F_{a^{\prime}}$. ∎
@@
Let $a,a^{\prime}\in I^{*}$ Then $B_{a^{\prime}}\subseteq^{*}B_{a}$ if and only if $a\leq a^{\prime}$.
@@
If $a=a^{\prime}$ then the implication is clear. If $a<a^{\prime}$ then $B_{a}\in F_{a^{\prime}}$ by Claim 4.4. Since $B_{a^{\prime}}$ is generic for the filter Mathias forcing defined from $F_{a^{\prime}}$, $B_{a^{\prime}}\subseteq^{*}B_{a}$.
Now let $a\not\leq a^{\prime}$. We work in the dense subset of $\mathbb{P}$ where $p\restriction I_{<b}$ decides the stem $s^{p}_{b}$ of the condition $p(b)$ for all $b\in\operatorname{dom}(p)$, and write $p(b)=(s^{p}_{b},\dot{A}^{p}_{b})$.
Assume for a contradiction that $p\in G$ and $p\Vdash B_{a^{\prime}}\setminus\zeta\subseteq B_{a}$ for some $\zeta<\kappa$. We may assume that $a,a^{\prime}\in\operatorname{dom}(p)$. Now let $I_{\leq a^{\prime}}=I_{<a^{\prime}}\cup\{a^{\prime}\}$, and note that $a\notin I_{\leq a^{\prime}}$. Extend $p\restriction I_{\leq a^{\prime}}$ to some $q\in{\mathbb{P}}_{I_{\leq a^{\prime}}}$ so that $\gamma\in s^{q}_{a^{\prime}}$ for some $\gamma>\zeta,\max(s^{p}_{a})$. Let $p_{1}=p\cup q$ and note that $p_{1}(a)=p(a)=(s^{p}_{a},\dot{A}^{p}_{a})$. Extend $p_{1}$ to $p_{2}$ such that $p_{2}(a)=(s^{p}_{a},\dot{A}^{p}_{a}\setminus(\gamma+1))$. Then $p_{2}\Vdash\mbox{``$\gamma\in\dot{B}_{a^{\prime}}\setminus\zeta$ and $\gamma\notin B_{a}$''}$ ” for an immediate contradiction. ∎
@@
In $V[G]$, $\{B_{a}:a\in I^{*}\}$ forms a basis for $U$.
@@
Let $Y\in V[G]$ with $Y\in U$. Then $Y\in V[G_{a}]$ for some $a\in I^{*}$. Arguing as in the proof of Claim 4.4, there is $b\in I^{*}$ such that $Y\in F_{b}$, so that $B_{b}\subseteq^{*}Y$. ∎
This concludes the proof of Theorem 4.2.
@@
Relative to a supercompact cardinal, it is consistent that there is a supercompact cardinal $\kappa$ such that for any regular cardinals $\kappa<\lambda_{0}<...<\lambda_{n}$, there is a $\kappa$-directed closed, $\kappa^{+}$-cc forcing extension where there is a normal ultrafilter over $\kappa$ which is a simple $P_{\lambda_{0}\times...\times\lambda_{n}}$-point.
Remark 4.8.
[…]
@@
    $|D|<\lambda=cf(\lambda)$.
@@
    $\mathbb{D}$ is $\kappa^{+}$-directed and well-founded.
The only modifications in the proof above are cosmetic:
  • work with the order $(d,\beta)\leq(d^{\prime},\beta^{\prime})$ if and only if $d\leq_{D}d^{\prime}\wedge\beta\leq\beta^{\prime}$.
  • replace $\prod_{0\leq i\leq n-1}\lambda_{i}$ by $\mathbb{D}$.
  • An ordinal $\beta<\lambda$ is active if $A_{\beta}$ guesses a $\mathbb{P}_{\mathbb{D}\times\beta}$-name for a measure. Note that $\mathbb{D}\times\beta$ consistes of pairs $(d,\gamma)$ such that $\gamma<\beta$. Consider a pair $a=(d,\beta)$, $F^{0}_{a}$ consists of $\mathbb{P}_{a}$-names (i.e. less or equal in both coordinates) $\dot{A}$ which are forced to be in $A_{\beta}$. Then $F^{1}_{a}$, $\mathcal{X}_{a}$ and $F_{a}$ are defined the same way,
  • extend the order on $\mathbb{D}\times\lambda$ to a well-ordering of ordertype $\lambda$.
  • use the $\kappa^{+}$-directedness of $\mathbb{D}\times\lambda$ to analyze as before how sets appear in the intermediate models, and the fact that $|\mathbb{D}|<\lambda$ to run the diamond arguments.
Corollary 4.9.
Relative to the existence of a supercompact cardinal $\kappa$, it is consistent that $\kappa$ is supercompact and for all regular cardinals $\kappa<\lambda_{0}<\lambda_{2}<...<\lambda_{n}$ there is a $\kappa^{+}$-cc and $\kappa$-closed poset forcing that there is a base $\mathcal{C}$ for the club filter $\text{Cub}_{\kappa}$ such that $(\mathcal{C},\supseteq^{*})\simeq\lambda_{0}\times\lambda_{2}\times...\times\lambda_{n}$.
@@
Let $\mathcal{B}$ be the base from the previous theorem and let $\mathcal{C}$ be the set of closures of the sets of $\mathcal{B}$. Then since $\mathcal{B}$ generated a normal ultrafilter, $\mathcal{C}$ generates $\text{Cub}_{\kappa}$. Clearly, if $X\subseteq^{*}Y$ then $cl(X)\subseteq^{*}cl(Y)$. A closer look at the proof of Claim 4.5 will reveal that if $X,Y\in\mathcal{B}$ and $X\not\subseteq^{*}Y$, then $X\not\subseteq^{*}cl(Y)$. Indeed, the choice of $\gamma$ is such that $\gamma$ is forced to be outside of the closure of $\dot{C}_{\alpha,\beta}$. Therefore, $cl(X)\not\subseteq^{*}cl(Y)$. ∎
5. Applications to cofinal types and cardinal characteristics
The following set is also known as the point-spectrum which was studied for example in [30, 25, 4, 24].
@@ 5. Applications to cofinal types and cardinal characteristics
Clearly, the Tukey spectrum is an invariant of the Tukey order. An equivalent condition for $\lambda\in\operatorname{Sp}_{T}(\mathbb{D})$ is to require the existence of a sequence ${\langle}d_{\alpha}\mid\alpha<\lambda{\rangle}$ such that for every $I\in[\lambda]^{\lambda}$, $\{d_{\alpha}\mid\alpha\in I\}$ is unbounded in $\mathbb{D}$. For the following fact, see for example [25]:
Fact 5.2.
@@
    If $\mathbb{P}$ is linear, then $\operatorname{Sp}_{T}(\mathbb{P})=\{cf(\mathbb{P})\}$.
@@
Suppose that $\mathbb{D}$ is a well-founded, degree $2$, countably directed semi-lattice, with no maximal elements, then there is a ccc forcing extension in which there is an ultrafilter $U$ on $\omega$ such that
$\operatorname{Sp}_{T}(U)=\operatorname{Sp}_{T}(\mathbb{D}).$
@@
Relative to the existence of a supercompact cardinal, there is a model $V$ with a supercompact cardinal $\kappa$ such that for every regular cardinals $\kappa<\lambda_{0}<\lambda_{1}<...<\lambda_{n}$, there is a ${<}\kappa$-directed closed $\kappa^{+}$-cc forcing extension in which there is a normal ultrafilter $U$ such that:
$\operatorname{Sp}_{T}(U)=\operatorname{Sp}_{T}(\lambda_{0}\times...\times\lambda_{n})=\{\lambda_{0},...,\lambda_{n}\}.$
In the previous corollary, $\operatorname{Sp}_{T}(\mathbb{D})$ is computed in the generic extension, which might be sensitive to forcing. The next proposition shows that no discrepancy arises in the interpretation of the spectrum when the generic extension is a ccc forcing extension. More specifically:
@@
[…]
@@
If $\lambda\in\operatorname{Sp}_{T}(\mathbb{D})^{V}$, then there is a sequence ${\langle}d_{\alpha}\mid\alpha<\lambda{\rangle}\subseteq D$ such that for any $I\in[\lambda]^{\lambda}$, $\{d_{\alpha}\mid\alpha\in I\}$ is unbounded in $\mathbb{D}$. To see that this sequence witnesses that $\lambda\in\operatorname{Sp}_{T}(\mathbb{D})^{V[G]}$, suppose not, then there is $I\in[\lambda]^{\lambda}$, $I\in V[G]$ such that $\{d_{\alpha}\mid\alpha\in I\}$ is bounded by some $d\in D$. Back in $V$, consider $I^{*}=\{\alpha<\lambda\mid d_{\alpha}\leq_{D}d\}$. Then $I\subseteq I^{*}$ and therefore $|I^{*}|=\lambda$. This is a contradiction since $\{d_{\alpha}\mid\alpha\in I^{*}\}$ should be unbounded.
For the converse, suppose that $\lambda\in\operatorname{Sp}_{T}(\mathbb{D})^{V[G]}$. Let $\langle d_{\alpha}\mid\alpha<\lambda\rangle\subseteq D$ be a sequence in $V[G]$ such that for each index set $I\in[\lambda]^{\lambda}$ the sequence $\langle d_{\alpha}\mid\alpha\in I\rangle$ is unbounded. Let $p\in G$ be a condition forcing the above. For each $\alpha<\lambda$ let $\dot{d}_{\alpha}$ be a $\mathbb{P}$-name such that $(\dot{d}_{\alpha})_{G}=d_{\alpha}$. For each $\alpha<\lambda$ let $p_{\alpha}\leq p$ such that $p_{\alpha}\Vdash_{\mathbb{P}}\dot{d}_{\alpha}=\check{e}_{\alpha}$ for some $e_{\alpha}\in D$. We claim that $\langle e_{\alpha}\mid\alpha<\lambda\rangle$ witnesses $\lambda\in\operatorname{Sp}(\mathbb{D})^{V}$. Suppose otherwise, and let $I\in[\lambda]^{\lambda}$ be such that $\langle e_{\alpha}\mid\alpha\in I\rangle$ is bounded. To contradict this assumption it suffices to check that $p\Vdash_{\mathbb{P}}``|\{\alpha<\lambda\mid p_{\alpha}\in\dot{G}\}|=\lambda"$. So, let us do it: Suppose towards a contradiction that this was false. Since $\mathbb{P}$ is $\lambda$-cc there is $\beta<\lambda$ such that $p\Vdash\{\alpha<\lambda\mid p_{\alpha}\in\dot{G}\}\subseteq\beta$. Note that this is impossible because $p_{\beta}$ itself forces ``$p_{\beta}\in\dot{G}"$. ∎
@@
[…]
@@
[…]
@@
and by Fact 5.2, $\operatorname{Sp}_{T}(U)=\{\lambda_{0},...\lambda_{n}\}$. By Remark 5.8 and Proposition 5.7, $\operatorname{Sp}_{dp}(\lambda_{0}\times...\times\lambda_{n})\subseteq\operatorname{Sp}_{dp}(U,\supseteq^{*})$. Finally note that for every $0\leq i\leq n$, $\lambda_{i}\in\operatorname{Sp}_{dp}(\lambda_{0}\times...\times\lambda_{n})$ as witnessed by the sequence
@@
    Let $\omega<\lambda_{0}<\lambda_{1}$ be regular cardinals. It is consistent that there is an ultrafilter on $\omega$, such that $\operatorname{Sp}_{dp}(U)=\operatorname{Sp}_{T}(U)=\{\lambda_{0},\lambda_{1}\}$.
@@
    Relative to a supercompact cardinal, it is consistent that there is supercompact cardinal $\kappa$ such that for any regular cardinals $\lambda_{0},\lambda_{1},...,\lambda_{n}>\kappa$, there is a $\kappa$-directed closed $\kappa^{+}$-cc forcing extension where:

### v2

@@ 4.1. At a measurable cardinal
Let $a=(\beta,d)$ and $a^{\prime}=(\beta^{\prime},d^{\prime})$. By the definition of the order, $\beta<\beta^{\prime}$. Then, clearly $B_{a}\in V[G_{\beta^{\prime}\times\mathbb{D}_{<a^{\prime}}}]$. Let $\dot{B}_{a}$ be a canonical ${\mathbb{P}}_{<a^{\prime}}$-name for $B_{a}$. We need to show that $\Vdash^{V}_{{\mathbb{P}}_{\beta^{\prime}\times\mathbb{D}}}\dot{B}_{a}\in A_{\beta^{\prime}}$. Let $G_{\beta^{\prime}\times\mathbb{D}}$ be an arbitrary generic for ${\mathbb{P}}_{\beta^{\prime}\times\mathbb{D}}$, and force to prolong it to a $\mathbb{P}$-generic $G$. Then by Claim 4.8 and Equation ($\star$4)
$B_{a}\in U\cap V[G_{\beta^{\prime}\times\mathbb{D}}]=\dot{U}_{\beta^{\prime}}[G_{\beta^{\prime}\times\mathbb{D}}]=A_{\beta^{\prime}}[G_{\beta^{\prime}\times\mathbb{D}}]$
Note that this is forced by the trivial condition. Hence, by definition, $(1,\dot{B}_{a})\in\dot{F}_{a^{\prime}}$ and $B_{a}\in F_{a^{\prime}}$. ∎
This concludes the proof of Theorem 4.7. ∎
We are ready to prove Theorem 4.6. Since ($\dagger$3.5) was already proven, we would like to apply Lemma 4.4 to some $I\subseteq S^{*}_{\dot{U}}$:
Proof of Theorem 4.6.
Let $\delta\in S^{*}_{\dot{U}}$ be a limit point of $S^{*}_{\dot{U}}$ such that there is $I\subseteq S^{*}_{\dot{U}}$ unbounded in $\delta$ with $otp(I)=\mu$. To find such an $\delta$, simply note that the set of all limit points of $S^{*}_{\dot{U}}$ which have such a set $I$ is a club relative to ${\rm cf}(\mu)$. Applying Lemma 4.4 or more precisely Remark 4.5, we conclude that $U\cap V[G_{\delta\times\mathbb{D}}]=\dot{U}_{\delta}[G_{\delta}]$ is a simple $\mathbb{P}_{\mathbb{D}}$-point. ∎
Corollary 4.10.
From the model of the previous theorem, for every $\kappa^{+}$-directed well-founded poset $\mathbb{D}$, there is a $\kappa^{+}$-cc and $\kappa$-directed closed forcing that adds a base $\mathcal{C}$ for the club filter $\text{Cub}_{\kappa}$ such that $(\mathcal{C},\supseteq^{*})\simeq\mathbb{D}$.
Proof.
Let $\mathcal{B}$ be the base from the previous theorem and let $\mathcal{C}$ be the set of closures of the sets of $\mathcal{B}$. Then since $\mathcal{B}$ generated a normal ultrafilter, $\mathcal{C}$ generates $\text{Cub}_{\kappa}$. Clearly, if $X\subseteq^{*}Y$ then $cl(X)\subseteq^{*}cl(Y)$. A closer look at the proof of Corollary 3.5 will reveal that if $X,Y\in\mathcal{B}$ and $X\not\subseteq^{*}Y$, then $X\not\subseteq^{*}cl(Y)$. Indeed, in Lemma 3.4, the choice of $\gamma$ is such that $\gamma$ is forced to be outside of the closure of $\dot{B}_{e}$. Therefore, $cl(X)\not\subseteq^{*}cl(Y)$. ∎
4.2. The countable case
Theorem 4.11.
Let $\mathbb{D}$ be a well-founded, $\sigma$-directed set, which has a possible order type $\lambda$, and assume that $\diamondsuit_{\lambda^{+}}(\lambda^{+}\cap\operatorname{cof}(\lambda))$ holds. Then, there is a ccc forcing extension in which there is a simple $P_{\mathbb{D}}$-point ultrafilter on $\omega$.
Proof.
Let $\mathbb{P}=\mathbb{P}_{\mathbb{D}^{+}}$ be the Mathias iteration from section §4.
Let us work in a generic extension $V[G]$. For each $a\in\mathbb{D}^{+}$, recall that $B_{a}$ denotes the Mathias real added by this coordinate.
Let $I\subseteq\lambda^{+}$ with the following properties:
  • $(A)$
    Every $\alpha\in I$ is active.
  • $(B)$
    For every $\alpha<\beta$ in $I$, for every $d\in\mathbb{D}$,
    $\Vdash_{\mathbb{P_{\beta\times\mathbb{D}}}}\dot{B}_{\alpha,d}\in A_{\beta}.$
    (Recall that $A_{\beta}$ denotes the $\beta^{\text{th}}$-element in the diamond sequence)
  • $(C)$
    II is maximal among sets with properties $(A),(B)$.
By Zorn’s lemma, $I$ exists, and it is defined in $V$. We will show that $I$ must be unbounded, and that $\{B_{(\alpha,d)}$: $\alpha\in I,d\in\mathbb{D}\}$ generates an ultrafilter.
@@ 4.2. The countable case
Let $G_{\alpha}\subseteq\mathbb{P}_{\alpha\times\mathbb{D}}$ be $V$-generic for some active level $\alpha$. Then the collection $\{B_{(\alpha,d)}$: $d\in\mathbb{D}\}$ generates a proper filter which extends $A_{\alpha}[G_{\alpha}]$.
@@
Let us show that the collection $\{B_{(\alpha,d)}\mid d\in\mathbb{D}\}$ has the finite intersection property. Assume otherwise, then working in $V[G_{\alpha}]$, there is a condition $p\in\mathbb{P}_{(\alpha+1)\times\mathbb{D}}$ with $p\restriction\alpha\times{\mathbb{D}}\in G_{\alpha}$ forcing that the intersection of some finite subset of $\{\dot{B}_{(\alpha,d)}\mid d\in\mathbb{D}\}$ is finite. By strengthening $p$ we may find a specific finite set $b\subseteq\mathbb{D}$ and $l<\omega$ such that $p$ forces $\bigcap_{x\in b}\dot{B}_{(\alpha,x)}\subseteq l$.
Let us look at $p(\alpha,x)=\langle\check{t}_{x},\dot{S}_{x}\rangle$. As we are working in $V[G_{\alpha}]$ we may treat $S_{x}=\dot{S}_{x}[G_{\alpha}]$ as actual sets and not merely names. As $S_{x}\in A_{\alpha}[G_{\alpha}]$ for each $x$, $\bigcap_{x\in b}S_{x}$ is infinite and in particular there is some $l\leq k<\omega$ in this intersection which is larger than $\max t_{x}$ for all $x\in b$. Extend the condition $p\restriction\alpha\times\mathbb{D}$ to $p^{*}\restriction\alpha\times\mathbb{D}\in G_{\alpha}$ so that $p^{*}\restriction\alpha\times\mathbb{D}\Vdash k\in\bigcap_{x\in b}\dot{S}_{x}$, and extend each $p(\alpha,x)$ to $p^{*}(\alpha,x)$ by adding $k$ to $t_{x}$ for all $x\in b$. $p^{*}$ is a legitimate condition extending $p$, which forces contradictory information.
The fact that the filter generated by $\{B_{(\alpha,d)}\mid d\in\mathbb{D}\}$ extends $A_{\alpha}[G_{\alpha}]$ follows from the observation in Remark 4.2(IIb). ∎
Corollary 4.13.
The collection $\{B_{(\alpha,a)}\mid\alpha\in I,a\in\mathbb{D}\}$ has the finite intersection property.
Proof.
We use Claim 4.12: Indeed, this would be trivial if $I$ happens to be empty, so suppose that $I\neq\emptyset$. If $I$ has a maximal element $\alpha_{*}$, and we force to obtain $G_{\alpha^{*}}$ which is ${\mathbb{P}}_{\alpha^{*}\times{\mathbb{D}}}$-generic, then $\{B_{(\alpha_{*},a)}\mid a\in\mathbb{D}\}\cup A_{\alpha^{*}}[G_{\alpha^{*}}]$ is a collection that contains $\{B_{(\beta,a)}\mid\beta\in I,a\in\mathbb{D}\}$ by property $(B)$, and by property $(A)$ we may apply Claim 4.12 to conclude that this collection has the finite intersection property.
Otherwise, for any given finite collection of elements in $I\times\mathbb{D}$, we can chop $I$ up to a sufficiently large ordinal $\alpha^{*}\in I$ and reduce the problem to the previous case.∎
Let $\mathcal{F}$ be the filter generated by $\{B_{(\alpha,d)}\mid\alpha\in I,d\in\mathbb{D}\}$ and let $\mathcal{U}$ be an ultrafilter extending it in $V[G]$. Let $\dot{U}$ be a nice name for the ultrafilter $\mathcal{U}$ with the property that every condition forces that it contains $\{B_{(\alpha,d)}\mid\alpha\in I,\,d\in\mathbb{D}\}$.
@@
II is unbounded, and $\{B_{(\alpha,d)}\mid\alpha\in I,\,d\in\mathbb{D}\}$ generates $\mathcal{U}$.
@@
For the first claim, assume that $I$ is bounded and let $\beta<\lambda^{+}$ large enough so that $I\subseteq\beta$. Pick $\gamma>\beta$ in which the restriction of $\dot{U}$ to $\mathbb{P}_{\gamma\times\mathbb{D}}$ is guessed by the diamond sequence. Let us denote this name for an ultrafilter by $\dot{U}_{\gamma}$ (i.e. $A_{\gamma}=\dot{U}_{\gamma}$). Clearly, $I\cup\{\gamma\}$ satisfies the properties (A)(B)(A)-(B), contradicting the maximality in property $(C)$.
For the second claim, since $\{B_{(\alpha,d)}\mid d\in\mathbb{D}\}$ generates in $V[G_{\alpha+1}]$ a filter extending the $V[G_{\alpha}]$-ultrafilter $A_{\alpha}[G_{\alpha}]$, we conclude that $\{B_{(\alpha,d)}\mid d\in\mathbb{D}\}$ generates a $V[G]$-ultrafilter which then has to be $\mathcal{U}$. ∎
To finish the proof, use Claim 4.14 to see that $I$ satisfies the assumption of Lemma 4.4 and produce the club $C$. Pick any point $\beta\in C\cap S^{*}_{\dot{U}}$, such that for some unbounded $I^{*}\subseteq I\cap\beta$, $\text{otp}(I^{*})=\lambda$. Such a point exists by the stationarity of $S^{*}_{\dot{U}}\cap C$ and the observation that there is a club relative to $\text{cof}(\lambda)$ of $\beta$’s such that there is $I^{*}\subseteq I$, $\text{otp}(I^{*})=\lambda$ and $I^{*}$ is unbounded in $\beta$. By the definition of $I$, $I^{*}\times\mathbb{D}$ satisfies ($\dagger$3.5) and by the previous claim $\{B_{(\alpha,d)}\mid\alpha\in I^{*},d\in\mathbb{D}\}$ generates an ultrafilter, so Lemma 4.4 applies. ∎
5. Applications to cofinal types and cardinal characteristics
In this section we derive some corollaries from our main result. Let us formulate a simple lemma which explains what happens to the isomorphism or Tukey type of the ordered set $\mathbb{D}$ under those mild forcing extensions. Clearly, changes are made (as we are adding an ultrafilter which is order isomorphic to it), but the essential structure of $\mathbb{D}$ is preserved. Note that in any forcing extension, a ground model unbounded (cofinal) map remains unbounded (cofinal).
Lemma 5.1.
Let $\mathbb{D}$ be a $\kappa^{+}$-directed set and $\mathbb{P}$ a $\kappa^{+}$-cc forcing and let $G\subseteq\mathbb{P}$ be $V$-generic. Then for any $f:A\to\mathbb{D}\in V[G]$, $A\in V$, there is $f^{\prime}:A\to\mathbb{D}\in V$ such that $f(a)\leq f^{\prime}(a)$ for all $a\in A$.
Proof.
This is straightforward from the $\kappa^{+}$-cc and $\mathbb{D}$ being $\kappa^{+}$-directed. ∎
@@ 5. Applications to cofinal types and cardinal characteristics
Let $A,\mathbb{D}\in V$ and assume that $\mathbb{D}$ is $\kappa^{+}$-directed, $\mathbb{P}$ is a $\kappa^{+}$-cc forcing notion and $G\subseteq\mathbb{P}$ is $V$-generic. Then:
@@
    $V[G]\models A\leq_{T}\mathbb{D}$ iff $V\models A\leq_{T}\mathbb{D}$.
@@
    $V[G]\models A\geq_{T}\mathbb{D}$ iff $V\models A\geq_{T}\mathbb{D}$.
@@
In both $(1)$ and $(2)$ we are left to show the implication from left to right. For $(1)$, let $f:A\to\mathbb{D}\in V[G]$ be unbounded and let $f^{\prime}:A\to\mathbb{D}\in V$ be the map from the previous lemma. We claim that $f^{\prime}$ is unbounded, otherwise, there is $\mathcal{A}\subseteq A$ unbounded and $d\in\mathbb{D}$ such that $f^{\prime}[\mathcal{A}]$ is bounded by $d$. Clearly, the boundedness of a subset of $\mathbb{D}$ is absolute, and since $f$ is an unbounded map, $f[\mathcal{A}]$ should be unbounded. But for every $a\in A$, $f(a)\leq f^{\prime}(a)\leq d$, contradiction. $(2)$ is proven similarly. ∎
The following set is also known as the point-spectrum. It was studied for example in [30, 24, 10, 23].
@@
Clearly, the Tukey spectrum is an invariant of the Tukey order and therefore of order isomorphism. An equivalent condition for $\lambda\in\operatorname{Sp}_{T}(\mathbb{D})$ is to require the existence of a sequence ${\langle}d_{\alpha}\mid\alpha<\lambda{\rangle}$ such that for every $I\in[\lambda]^{\lambda}$, $\{d_{\alpha}\mid\alpha\in I\}$ is unbounded in $\mathbb{D}$. Corollary 5.2 has the following corollary:
Corollary 5.4.
Let $\mathbb{P}$ be either $\kappa^{+}$-cc or $|\mathbb{D}|^{+}$-closed forcing and $\mathbb{D}$ a $\kappa^{+}$-directed set. Let $G\subseteq\mathbb{P}$ be a $V$-generic filter, then
$\operatorname{Sp}_{T}(\mathbb{D})^{V}=\operatorname{Sp}_{T}(\mathbb{D})^{V[G]}.$
Proof.
Note that $\operatorname{Sp}_{T}(\mathbb{D})\subseteq[\kappa^{+},|\mathbb{D}|]$, hence Corollary 5.2 implies the result for $\kappa^{+}$-cc forcings instantly. If $\mathbb{P}$ is $|\mathbb{D}|^{+}$-closed then this is much simpler since all the witnessing sequences for $\lambda$’s being in the Tukey spectrum exist in the ground model. ∎
For the following fact, see for example [24]:
Fact 5.5.
@@
    If $\mathbb{P}$ is linear, then $\operatorname{Sp}_{T}(\mathbb{P})=\{{\rm cf}(\mathbb{P})\}$.
@@
Suppose that $\mathbb{D}$ is a well-founded, countably directed set with $|\mathbb{D}|=\mu$. Then there is a ($\mu^{+}$-closed $\ast$ ccc) forcing extension in which there is an ultrafilter $U$ on $\omega$ such that $U\equiv_{T}\mathbb{D}$ and in particular $\operatorname{Sp}_{T}(U)=\operatorname{Sp}_{T}(\mathbb{D}).$
@@
Suppose that $\kappa$ is a supercompact cardinal and let $\mathbb{D}$ be a $\kappa^{+}$-directed well-founded set with $|\mathbb{D}|=\mu$. Then there is a ($\kappa$-$c$ $\ast\ \mu^{+}$-closed $\ast\ \kappa^{+}$-cc) forcing extension in which there is a normal ultrafilter $U$ on $\kappa$ such that $U\equiv_{T}\mathbb{D}$, and in particular $\operatorname{Sp}_{T}(U)=\operatorname{Sp}_{T}(\mathbb{D}).$
As for every normal ultrafilter $U$, the order $(U,\supseteq^{*})$ is $\kappa^{+}$-directed, this result is locally optimal — every partial order that might be compatible with the order $\supseteq^{*}$ of a normal ultrafilter can consistency be forced to actually be Tukey equivalent to one.
The next proposition shows that for the point spectrum to be preserved, we do not need any directedness assumptions. More specifically:
@@
Let $\mathbb{P}$ be a $\lambda$-cc forcing and $\mathbb{D}$ a directed set. Let $G\subseteq\mathbb{P}$ be a $V$-generic filter, then
@@
The implication from left to right always hold true. For the converse, suppose that $\lambda\in\operatorname{Sp}_{T}(\mathbb{D})^{V[G]}$. Let $\langle d_{\alpha}\mid\alpha<\lambda\rangle\subseteq D$ be a sequence in $V[G]$ such that for each index set $I\in[\lambda]^{\lambda}$ the sequence $\langle d_{\alpha}\mid\alpha\in I\rangle$ is unbounded. Let $p\in G$ be a condition forcing the above. For each $\alpha<\lambda$ let $\dot{d}_{\alpha}$ be a $\mathbb{P}$-name such that $(\dot{d}_{\alpha})_{G}=d_{\alpha}$. For each $\alpha<\lambda$ let $p_{\alpha}\leq p$ such that $p_{\alpha}\Vdash_{\mathbb{P}}\dot{d}_{\alpha}=\check{e}_{\alpha}$ for some $e_{\alpha}\in D$. We claim that $\langle e_{\alpha}\mid\alpha<\lambda\rangle$ witnesses $\lambda\in\operatorname{Sp}(\mathbb{D})^{V}$. Suppose otherwise, and let $I\in[\lambda]^{\lambda}$ be such that $\langle e_{\alpha}\mid\alpha\in I\rangle$ is bounded. To contradict this assumption it suffices to check that $p\Vdash_{\mathbb{P}}``|\{\alpha<\lambda\mid p_{\alpha}\in\dot{G}\}|=\lambda$". So, let us do it: Suppose towards a contradiction that this was false. Since $\mathbb{P}$ is $\lambda$-cc there is $\beta<\lambda$ such that $p\Vdash\{\alpha<\lambda\mid p_{\alpha}\in\dot{G}\}\subseteq\beta$. Note that this is impossible because $p_{\beta}$ itself forces $p_{\beta}\in\dot{G}$". ∎
@@
Let $\mathbb{D}=(D,\leq_{D})$ be a directed set. A $\mathbb{D}$-tower of length $\lambda$ is a sequence ${\langle}d_{\alpha}\mid\alpha<\lambda{\rangle}$ which is $\leq_{D}$-increasing and unbounded in $\mathbb{D}$. The depth spectrum of $\mathbb{D}$, denoted by $\operatorname{Sp}_{dp}(\mathbb{D})$ consists of all regular cardinals $\lambda$ such that there is a $\mathbb{D}$-tower of length $\lambda$.
@@
If $\mathcal{B}$ is cofinal in $\mathbb{D}=(D,\leq_{D})$, then [^2] Formally, $\mathbb{D}\equiv_{T}(\mathcal{B},\leq_{D}\restriction\mathcal{B})$., $\mathbb{D}\equiv_{T}\mathcal{B}$. Moreover, the identity map from $\mathcal{B}$ into $D$ is a monotone Tukey reduction.
@@
and by Fact 5.5, $\operatorname{Sp}_{T}(U)=\{\lambda_{0},...\lambda_{n}\}$. By Remark 5.11 and Proposition 5.10, $\operatorname{Sp}_{dp}(\lambda_{0}\times...\times\lambda_{n})\subseteq\operatorname{Sp}_{dp}(U,\supseteq^{*})$. Finally note that for every $0\leq i\leq n$, $\lambda_{i}\in\operatorname{Sp}_{dp}(\lambda_{0}\times...\times\lambda_{n})$ as witnessed by the sequence
@@
    Let $\omega<\lambda_{0}<\lambda_{1}<\lambda_{2}<\cdots<\lambda_{n}$ be regular cardinals. It is consistent that there is an ultrafilter on $\omega$, such that $\operatorname{Sp}_{dp}(U)=\operatorname{Sp}_{T}(U)=\{\lambda_{0},\lambda_{1},\lambda_{2},\dots,\lambda_{n}\}$.
@@
    Suppose that $\kappa$ is supercompact. Then for any regular cardinals $\kappa<\lambda_{0}<\lambda_{1}<...<\lambda_{n}$, there is a ($\kappa$-$c$ $\ast\ \lambda_{n}^{+}$-closed $\ast\ \kappa^{+}$-cc) extension where: